OpenStudy (anonymous):

Part 1: Write the equation of the line that passes through the points (–1, –4) and (2, 5) in point-slope form. Part 2: Using complete sentences, explain whether or not it matters which point is used in the final answer. Also explain why you chose the point you did.

OpenStudy (math2400):

Let me redeem my self haha.. Part 1: find the slope by using  so you'll get:  now that you have your slope plug in your values to the point slope formula using your first given point: y+4 = 3(x+1) which simplifies to y = 3x-1